OpenStudy (anonymous):

Will Fan And Medal!!!!!! Lori draws a clear marble from a bag of marbles 14 times out of a toal of 56 drwas with replacement. If there are 16 marbles in the bag in all, about how many of the marbles are clear? A.4 B.8 C.2 D.12

OpenStudy (kirbykirby):

Well you might expect that since she drew a marble 14 times out of 56, this is \(\Large \frac{14}{56}=\frac{1}{4}\) . So a quarter of the time, she picks a clear marble. So if there are 16 marbles, you expect 1/4 of them to be clear.. thus \(\large 16 \times \frac{1}{4}=4\). This is of course assuming that her draws are uniformly randomly distributed.

OpenStudy (anonymous):

Thanks that helped a lot:) @kirbykirby